solve: (3/n)-(7/square root of n)-6 =0 Thank you (:
Answer by stanbon(75887)

solve: (3/n)-(7/sqrt(n))-6 =0
Multiply through by "n" to get:
3-7sqrt(n)=6n
Then -7sqrt(n)=6n-3
Square both sides to get:
49n=36n^2-36n+9
36n^2-85n+9=0
(4n-9)(9n-1)=0
n=9/4 or n=1/9
Checking you will find that n=9/4 is extraneous.
Hope this helps.
Cheers,
Stan H.